OpenStudy (anonymous):

how do i solve this: twice as many dimes as nickels; value of dimes is .75 more than value of nickels. How many nickels?

OpenStudy (anonymous):

You want to set up two equations. One giving value to the coins, the other accounting for the number of coins. d=dimes n=nickels So first, the number of each: since you have twice as many dimes as nickels, if you multiply your number of nickels by 2, you'll get your number of dimes. So 2n=d. Your second equation would deal with value. A dime is worth 10 cents, so 10d=value of a dime and 5n=value of a nickel. Value of nickels + .75 =value of dimes. So 10d=5n+75. Now solve your two equations by substituting 2n for d in your second equation since 2n=d. So 10d=5n+75 10(2n)=5n+75 20n=5n+75 15n=75 n=5 Now that you know you have 5 nickels, plug into your equation and solve for d. 2n=d 2(5)=d d=10 So you have 10 dimes and 5 nickels.

OpenStudy (anonymous):

thx!